OpenStudy (anonymous):

Please help, how would you simplify 8^4/3 I'm thinking it's 16.

OpenStudy (saifoo.khan):

your thinking is correct.

OpenStudy (anonymous):

no its not its 1365.3

OpenStudy (saifoo.khan):

Wrong^

OpenStudy (anonymous):

then how do you explaine

OpenStudy (anonymous):

.......................................................................

OpenStudy (saifoo.khan):

Check, 18 * 18 * 18 * 18 = 4096 16 * 16 * 16 = 4096.

OpenStudy (anonymous):

8 not 18

OpenStudy (anonymous):

u messed up

OpenStudy (saifoo.khan):

\[\Huge \sqrt[3]{8^4}\]\[\Huge \sqrt[3]{4096}\]\[\Huge 16\]

OpenStudy (anonymous):

lorda mercy. the cube root of 8 is 2. 2 to the power of 4 is 16 halas

OpenStudy (saifoo.khan):

Check, 8 * 8 * 8 * 8 = 4096 16 * 16 * 16 = 4096.

OpenStudy (saifoo.khan):

Typos!

OpenStudy (saifoo.khan):

lol @sat

OpenStudy (anonymous):

it may be 4096 but divide that by 3 you get 1365.3

OpenStudy (anonymous):

so ha

OpenStudy (anonymous):

a speachless dude

OpenStudy (anonymous):

..........................